# Enhancing Morphological Measurements of Cosmic Web with Delaunay Tessellation Field Estimation

### v1

@@ 5.2 The MFs with Gaussian smoothing
Figure 4: The DTFE MFs of uniform-weighted halo fields with various grid resolutions in real space. The corresponding halo number density for these fields is $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$. The blue, red, and green lines correspond to halo fields with $N_{\mathrm{g}}=1024^{3}$, $N_{\mathrm{g}}=2048^{3}$, and $N_{\mathrm{g}}=4096^{3}$ grid cells, respectively. As shown, the amplitudes of $V_{1}$, $V_{2}$, and $V_{3}$ become greater as $N_{\mathrm{g}}$ increases. This means that, under DTFE scheme, higher grid resolution can help resolve more subtle structural elements from point set.
@@ 6 Systematic Effects on DTFE MFs of Halo Fields
Using a state-of-the-art simulation (i.e., TianZero; cf. Section 4), in this section, we investigate various dominant systematic effects on DTFE MFs, which are caused by finite voxel sizes, halo number densities, halo weighting schemes, and RSDs, with the goal to comprehend how these effects alter the morphological measurements with DTFE MFs. Note that, starting from this section, we perform the analyses for DTFE MFs at halo field level.
@@ 6.2 Systematic effects from halo number densities
6.3 Systematic effects from halo weighting schemes
DTFE halo fields can be constructed under different halo weightings (cf. Section 3), which can yield different halo biases (Liu et al. 2021). In this work, we consider three weighting schemes: uniform, mass (Seljak et al. 2009), and optimal [^10] The ‘optimal weighting’ refers to the weighting scheme that can minimize the stochasticity of halos with respect to underlying dark matter (Hamaus et al. 2010; Cai et al. 2011; Liu et al. 2021). (Hamaus et al. 2010; Cai et al. 2011) weightings, represented by the forms $w_{i}(M)=1,M_{i},M_{i}+M_{0}$, respectively. Here, $M_{0}$ is a free parameter, and we adopt $M_{0}=3M_{\mathrm{min}}$ [^11] Note that this is an empirical relation fund in Hamaus et al. 2010, which may not hold under a different condition, e.g., using a different halo-finder algorithm (cf. Liu et al. 2021). Here, we just naively employ this relation as a proxy for the optimal weighting in our tests. (Hamaus et al. 2010), where $M_{\mathrm{min}}$ is the mass cutoff (cf. Section 4). Since the last two weightings both depend on halo masses, we collectively denote them as mass-dependent weightings. These two weightings have been utilized in various cosmological studies [e.g., primordial non-Gaussianities (Hamaus et al. 2011), growth rate of structure formation (Hamaus et al. 2012), and initial condition reconstruction (Liu et al. 2021), etc.], given that they can significantly improve the correlation between halo field and underlying matter field (cf. Seljak et al. 2009; Hamaus et al. 2010; Liu et al. 2021).
Compared to uniform weighting, mass-dependent weightings trend to upweight the regions with higher halo abundance, as these regions are more likely to contain massive halos (cf. Haas et al. 2012; Zhao et al. 2015). As a result, the density contrasts between high- and low-density regions are intensified in mass-dependent weighted halo fields, leading to the stretching of corresponding MFs in two opposite directions (cf. Fig. 6 and Fig. 7). Also, due to halo fields are dominated by low-density regions, the MFs of mass-dependent weighted halo fields are visually shifted towards the direction of low-density thresholds. Moreover, we observe that the amplitudes of $V_{1}$, $V_{2}$, and $V_{3}$ for mass-dependent weightings are enhanced relative to the case of uniform weighting. This implies that mass-dependent weightings presumably aid in resolving more structural elements when estimating halo fields, consistent with previous findings that mass information can enhance the correlation between halo field and underlying matter field (Seljak et al. 2009; Hamaus et al. 2010; Liu et al. 2021).
@@ 6.4 Systematic effects from RSDs
Since valuable information about velocity field is encoded in RSD effects, an accurate modelling for these effects on MFs would open up a unique window to extract information on structure growth rate. In linear regime, Matsubara 1996 fund that Kaiser’s effect on genus (equivalent to $V_{3}$) can be predicted by
@@
DTFE MFs can naturally capture non-linear structure formation signatures induced by FOG effect. This is because DTFE tends to preserve maximum amount of non-linear structures, which are typically smoothed out in the CIC+GS scheme with a large $R_{\mathrm{G}}$ (cf. Kim et al. 2014 and Appleby et al. 2021b for (quasi-) linear-scale scenarios). Therefore, our results do not fit well with linear theory predictions (e.g., the Equation (15) under Kaiser approximation; also cf. Fig. 8). Indeed, using $N$-body simulations, Matsubara & Suto 1996 found that the amplitude of genus is more suppressed than expected by linear theory (also cf. Kim et al. 2014), consistent with our results (cf. the bottom-right panel of Fig. 8). Nevertheless, in principle, we can still apply emulator-based approaches to parameter estimates (cf. Marques et al. 2019; Petri et al. 2015), where significant RSD signatures should be critical to improving constraining power of MFs on cosmological parameters. This subject is beyond the scope of this work, and we defer the investigations to future studies.
@@ 7.1 The discriminative power for RSDs
Figure 10: The correlation matrices for DTFE MFs of uniform-weighted halo fields with different number densities. For left to right, each panel shows the case of halo number density of $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, $1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$, and $1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$, respectively.
Figure 11: The fisher error forecasts for structure growth rate, with $n_{\mathrm{b}}=32$ (cf. Section C for technical details). The results for CIC+GS and DTFE MFs are marked with blue circles and red diamonds, respectively. As shown, the y-axis values of red diamonds are much lower than those of blue circles by a factor of around $3$-$5$. This means that, compared with traditional CIC+GS MFs, DTFE MFs are more sensitive to cosmological parameters, because they contain more morphological information of LSS.
@@
We see that, for $V_{1}$, $V_{2}$, and $V_{3}$, RSD signals in DTFE MFs are significantly higher than those in CIC+GS MFs, regardless of halo number density. And, our results are well consistent with those found in previous works (e.g., Kim et al. 2014; Jiang et al. 2021; Lippich & Sánchez 2021; Appleby et al. 2021b), but showing more pronounced signals in DTFE case. Moreover, intriguingly, despite our analyses being performed at halo field level, we find that the signs and trends of $\Delta V_{i}\left(\nu\right)$ in our work are broadly the same with those in Jiang et al. 2021, where CIC+GS DM fields were employed for analyses. For interpretations of $\Delta V_{i}\left(\nu\right)$, we refer interested readers to that paper to find more details. In Fig. 13, we also present the S/N ratios of RSD effects, defined as $\left|\Delta V_{i}\right|/\sigma_{i}$. As shown, the amplitudes of RSD S/N ratios for DTFE MFs are also larger than those for CIC+GS MFs, except for $\left|\Delta\mathrm{V}_{0}\right|/\sigma_{0}$ with $\bar{n}_{\mathrm{h}}=1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$ (cf. the bottom subplot in left-top panel of Fig. 13), where the amplitudes in the two cases are basically comparable.
@@ 8 Summaries and Discussions
  • At halo field level, we first numerically study various dominant systematic effects on DTFE MFs, induced by finite voxel size, halo number densities, halo weighting schemes, and redshift space distortions. Then, we showcase the robust statistical power of DTFE MFs for extracting cosmological information encoded in RSDs. We find that DTFE MFs remarkably outperform traditional CIC+GS MFs by $\sim$ $2$ orders of magnitude in discriminative power for RSD effects and by a factor of $\sim$ $3$-$5$ in constraining power on structure growth rate. This is because that DTFE scheme can help conserve maximum morphological information of LSS from sparse tracers, rendering DTFE MFs more sensitive to cosmological parameters.
In view of the strong statistical power of DTFE MFs, we will employ this method to extract various critical cosmological signatures [e.g., neutrino masses (Liu et al. 2020), modified gravities (Fang et al. 2017), and primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008] imprinted on halo/galaxies fields, in our ongoing projects.
The implementation scheme for measuring MFs proposed in this paper is conceptually related to other Delaunay-based methods developed in previous works (Aragon-Calvo et al. 2010; Lippich & Sánchez 2021). These works compute MFs from triangulated iso-density surfaces, which are directly specified from Delaunay tessellation of a point set, without interpolating density field onto a regular grid. When density values at tessellation vertexes in these methods are estimated in the same manner as DTFE, their performance in measuring MFs should be highly similar to that of our method, with an added advantage of being free from finite voxel size issue. Alternatively, there might also be a possibility to directly evaluate MFs from alpha shapes of discrete tracers, which are specified via filtrations of Delaunay tessellation (van de Weygaert et al. 2011). Despite all that, the method presented in this work stands out by its simplicity and convenience in implementation, as it does not require any additional complex code designs for extracting MFs from triangulated isosurfaces. This is undeniably one of the main merits of our method.
@@ Acknowledgments
Y.L. would like to thank Cheng Zhao and Charling Tao for useful communications. We also thank the anonymous referee for their very detailed and helpful comments. This work was supported by National Science Foundation of China (No. 12303005, 12273015, 12273020, 11621303, 11890691), National Key R&D Program of China (No. 2023YFA1605601, 2023YFA1607800, 2023YFA1607802), and the science research grants from China Manned Space Project with No. CMS-CSST-2021-A03. Y.L. acknowledges supports from Shuimu Tsinghua Scholar Program (No. 2022SM173). This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
@@ Appendix A The Determinations of Smoothing Lengths
In this appendix, we illustrate our strategy for determining the smoothing lengths adopted in Section 5.2 and display the effects of shot noise on CIC+GS MFs. In previous works (e.g., Park et al. 2005; Choi et al. 2010; Parihar et al. 2014; Kim et al. 2014), smoothing lengths were empirically determined with the restriction $R_{G}\geq\bar{d}\sim\bar{n}_{\mathrm{tracer}}^{-1/3}$, where $\bar{n}_{\mathrm{tracer}}$ is tracer number density. Actually, in our tests, setting $R_{G}=\bar{d}$ cannot provide adequate smoothing for completely eliminating shot noise effects. Our strategy is to strike a balance between removing shot noise effects and retaining sufficient information. To achieve this, $R_{G}$ is determined through an iterative and step-nice refined process, such that the differences of MFs between downgraded field and noise-free field are controlled within $1$-$\sigma$ deviation. We note that this strategy seems to have never been proposed before. Therefore, it is worth exploring it theoretically.
@@ Appendix C The Methodology of Fisher Forecasts for Errors on Structure Growth Rate
Figure 14: The MFs of subfields. The upper panels are for CIC+MFs, and the lower panels are for DTFE MFs. For left to right, each panel shows the case of halo number density of $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, $1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$, and $1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$, respectively. For each panel, we show 100 sets of MFs, which are randomly selected from the MFs of 512 subfields. Here, each $V_{i}$ has $n_{\mathrm{b}}=64$ data points, and the unshielded regions indicate the density threshold ranges actually used in our calculations for Fisher forecasts (cf. Table 1).
Figure 15: The same as Fig. 11, but for the case of $n_{\mathrm{b}}=64$.
@@
In our scenario, data vectors refer to the MFs in redshift space $V_{i}^{(\mathrm{z})}(\nu)$, and data realizations refer to the $V_{i}^{(\mathrm{z})}(\nu)$ measured from $n_{\mathrm{f}}=8^{3}=512$ subfields (cf. Section 2). Specifically, the covariance matrices are first calculated by using the $512$ subfields with RSDs and then are rescaled by $1/n_{\mathrm{f}}$ [^14] This method is based on an implicit assumption that all subfields are statistically independent. This is not really satisfied in practice. Therefore, it only gives rough estimations (cf. Lacasa & Kunz 2017; Jiang et al. 2023). Since we mainly focus on comparisons of the two schemes for measuring MFs, this issue, in principle, does not affect our conclusions. To solve this issue, one can use a large number of simulation realizations to give accurate but expensive covariance estimates. We leave this for future works.. To calculate the derivatives in Equation C1, we construct two additional halo fields with modified RSDs by artificially increasing and decreasing the peculiar velocity of each halo by a factor of $0.03$, respectively. Then, the response functions are given as $\Delta\boldsymbol{\mu}=\boldsymbol{\mu}_{\mathrm{v}_{+}}-\boldsymbol{\mu}_{\mathrm{v}_{-}}=\Delta V_{i}^{(\mathrm{z})}(\nu)$, where subscript $\mathrm{v}_{+}$ and $\mathrm{v}_{-}$ indicate the two cases of increasing and decreasing halo peculiar velocities, respectively. Moreover, in RSD studies, there is an approximate relational expression between structure growth rate $f\sigma_{8}(z)$ and halo/galaxy velocity bias $b_{\mathrm{v}}$ (cf. Chen et al. 2018),
@@
In our calculation, we select $n_{\mathrm{b}}$ data points for each $V_{i}(\nu)$. These data points are evenly distributed within $[\nu_{\mathrm{min}},\nu_{\mathrm{max}}]$ in logarithmic coordinate. To avoid any problems sourced from the irreversibility of covariance matrices, the used density threshold range $[\nu_{\mathrm{lower}},\nu_{\mathrm{upper}}]$ should be narrowed as halo number density decreases. It will result in a decrease in the number of data points actually used in calculations. We illustrate this in Fig. 14151515We only display the results of $n_{\mathrm{b}}=64$, considering the striking similarity between these results and those of $n_{\mathrm{b}}=32$. (cf. the unshielded regions) and Table 1. In order to draw conclusions as reliable as possible, two $n_{\mathrm{b}}$ cases are tested, namely $n_{\mathrm{b}}=32$ (cf. Fig. 11) and $n_{\mathrm{b}}=64$ (cf. Fig. 15). This provides a consistency check for our methodology. In general, in comparison with Fig. 11, the predictive errors in Fig. 15 are relatively smaller, depending on different cases. Specifically, as the number of data points increases, we find the followings:
@@
    For DTFE MFs, the predictive errors remain relatively stable in the case of $n_{\mathrm{h}}=1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, but showing a noticeable decrease in the cases of $n_{\mathrm{h}}=1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$ and $n_{\mathrm{h}}=1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$.

### v2

@@ 5.2 The MFs with Gaussian smoothing
Figure 4: The DTFE MFs of uniform-weighted halo fields with various grid resolutions in real space. The corresponding halo number density for these fields is $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$. The blue, red, and green lines correspond to halo fields with $N_{\mathrm{g}}=1024^{3}$, $N_{\mathrm{g}}=2048^{3}$, and $N_{\mathrm{g}}=4096^{3}$ grid cells, respectively. As shown, the amplitudes of $V_{1}$, $V_{2}$, and $V_{3}$ become greater as $N_{\mathrm{g}}$ increases. This means that under DTFE scheme, higher grid resolution can help resolve more subtle structural elements from point set.
@@ 6 Systematic Effects on DTFE MFs of Halo Fields
Using a state-of-the-art simulation (i.e., TianZero; cf. Section 4), in this section, we investigate various dominant systematic effects on DTFE MFs, which are caused by finite voxel sizes, halo number densities, halo-weighting schemes, and RSDs, with the goal to comprehend how these effects alter the morphological measurements with DTFE MFs. Note that, starting from this section, we perform the analyses for DTFE MFs at halo field level.
@@ 6.2 Systematic effects from halo number densities
6.3 Systematic effects from halo-weighting schemes
DTFE halo fields can be constructed under different halo weightings (cf. Section 3), which can yield different halo biases (Liu et al. 2021). In this work, we consider three weighting schemes: uniform, mass (Seljak et al. 2009), and optimal [^10] The ‘optimal weighting’ refers to the weighting scheme that can minimize the stochasticity of halos with respect to underlying dark matter (Hamaus et al. 2010; Cai et al. 2011; Liu et al. 2021). (Hamaus et al. 2010; Cai et al. 2011) weightings, represented by the forms $w_{i}(M)=1,M_{i},M_{i}+M_{0}$, respectively. Here, $M_{0}$ is a free parameter, and we adopt $M_{0}=3M_{\mathrm{min}}$ [^11] Note that this is an empirical relation found in Hamaus et al. 2010, which may not hold under a different condition, e.g., using a different halo-finder algorithm (cf. Liu et al. 2021). Here, we just naively employ this relation as a proxy for the optimal weighting in our tests. (Hamaus et al. 2010), where $M_{\mathrm{min}}$ is the mass cutoff (cf. Section 4). Since the last two weightings both depend on halo masses, we collectively denote them as mass-dependent weightings. These two weightings have been utilized in various cosmological studies [e.g., primordial non-Gaussianities (Hamaus et al. 2011), growth rate of structure formation (Hamaus et al. 2012), and initial condition reconstruction (Liu et al. 2021), etc.], given that they can significantly improve the correlation between halo field and underlying matter field (cf. Seljak et al. 2009; Hamaus et al. 2010; Liu et al. 2021).
Compared to uniform weighting, mass-dependent weightings tend to upweight the regions with higher halo abundance, as these regions are more likely to contain massive halos (cf. Haas et al. 2012; Zhao et al. 2015). As a result, the density contrasts between high- and low-density regions are intensified in mass-dependent weighted halo fields, leading to the stretching of corresponding MFs in two opposite directions (cf. Fig. 6 and Fig. 7). Also, due to halo fields being dominated by low-density regions, the MFs of mass-dependent weighted halo fields are visually shifted towards the direction of low-density thresholds. Moreover, we observe that the amplitudes of $V_{1}$, $V_{2}$, and $V_{3}$ for mass-dependent weightings are enhanced relative to the case of uniform weighting. This implies that mass-dependent weightings presumably aid in resolving more structural elements when estimating halo fields, consistent with previous findings that mass information can enhance the correlation between halo field and underlying matter field (Seljak et al. 2009; Hamaus et al. 2010; Liu et al. 2021).
@@ 6.4 Systematic effects from RSDs
Since valuable information about velocity field is encoded in RSD effects, an accurate modelling for these effects on MFs would open up a unique window to extract information on structure growth rate. In linear regime, Matsubara 1996 found that Kaiser’s effect on genus (equivalent to $V_{3}$) can be predicted by
@@
DTFE MFs can naturally capture non-linear structure formation signatures induced by FOG effect. This is because DTFE tends to preserve maximum amount of non-linear structures, which are typically smoothed out in the CIC+GS scheme with a large $R_{\mathrm{G}}$ (cf. Kim et al. 2014 and Appleby et al. 2022 for (quasi-) linear-scale scenarios). Therefore, our results do not fit well with linear theory predictions (e.g., the Equation (15) under Kaiser approximation; also cf. Fig. 8). Indeed, using $N$-body simulations, Matsubara & Suto 1996 found that the amplitude of genus is more suppressed than expected by linear theory (also cf. Kim et al. 2014), consistent with our results (cf. the bottom-right panel of Fig. 8). Nevertheless, in principle, we can still apply emulator-based approaches to parameter estimates (cf. Marques et al. 2019; Petri et al. 2015), where significant RSD signatures should be critical to improving constraining power of MFs on cosmological parameters. This subject is beyond the scope of this work, and we defer the investigations to future studies.
@@ 7.1 The discriminative power for RSDs
Figure 10: The correlation matrices for DTFE MFs of uniform-weighted halo fields with different number densities. From left to right, each panel shows the case of halo number density of $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, $1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$, and $1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$, respectively.
Figure 11: The Fisher error forecasts for structure growth rate, with $n_{\mathrm{b}}=32$ (cf. Section C for technical details). The results for CIC+GS and DTFE MFs are marked with blue circles and red diamonds, respectively. As shown, the y-axis values of red diamonds are much lower than those of blue circles by a factor of around $3$-$5$. This means that compared with traditional CIC+GS MFs, DTFE MFs are more sensitive to cosmological parameters, because they contain more morphological information of LSS.
@@
We see that, for $V_{1}$, $V_{2}$, and $V_{3}$, RSD signals in DTFE MFs are significantly higher than those in CIC+GS MFs, regardless of halo number density. And, our results are well consistent with those found in previous works (e.g., Kim et al. 2014; Jiang et al. 2021; Lippich & Sánchez 2021; Appleby et al. 2022), but showing more pronounced signals in DTFE case. Moreover, intriguingly, despite our analyses being performed at halo field level, we find that the signs and trends of $\Delta V_{i}\left(\nu\right)$ in our work are broadly the same as those in Jiang et al. 2021, where CIC+GS DM fields were employed for analyses. For interpretations of $\Delta V_{i}\left(\nu\right)$, we refer interested readers to that paper to find more details. In Fig. 13, we also present the S/N ratios of RSD effects, defined as $\left|\Delta V_{i}\right|/\sigma_{i}$. As shown, the amplitudes of RSD S/N ratios for DTFE MFs are also larger than those for CIC+GS MFs, except for $\left|\Delta\mathrm{V}_{0}\right|/\sigma_{0}$ with $\bar{n}_{\mathrm{h}}=1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$ (cf. the bottom subplot in left-top panel of Fig. 13), where the amplitudes in the two cases are basically comparable.
@@ 8 Summaries and Discussions
  • At halo field level, we first numerically study various dominant systematic effects on DTFE MFs, induced by finite voxel sizes, halo number densities, halo-weighting schemes, and redshift space distortions. Then, we showcase the robust statistical power of DTFE MFs for extracting cosmological information encoded in RSDs. We find that DTFE MFs remarkably outperform traditional CIC+GS MFs by $\sim$ $2$ orders of magnitude in discriminative power for RSD effects and by a factor of $\sim$ $3$-$5$ in constraining power on structure growth rate. This is because DTFE scheme can help conserve maximum morphological information of LSS from sparse tracers, rendering DTFE MFs more sensitive to cosmological parameters.
In view of the strong statistical power of DTFE MFs, we will employ this method to extract various critical cosmological signatures [e.g., neutrino masses (Liu et al. 2020), modified gravities (Fang et al. 2017), and primordial non-Gaussianities (Hikage et al. 2006; Hikage et al. 2008)] imprinted on halo/galaxy fields, in our ongoing projects.
The implementation scheme for measuring MFs proposed in this paper is conceptually related to other Delaunay-based methods developed in previous works (Aragon-Calvo et al. 2010; Lippich & Sánchez 2021). These works compute MFs from triangulated iso-density surfaces, which are directly specified from Delaunay tessellation of a point set, without interpolating density field onto a regular grid. When density values at tessellation vertexes in these methods are estimated in the same manner as DTFE, their performance in measuring MFs should be highly similar to that of our method, with an added advantage of being free from finite voxel size issue. Alternatively, there might also be a possibility to directly evaluate MFs from alpha shapes of discrete tracers, which are specified via filtrations of Delaunay tessellation (van de Weygaert et al. 2011). Despite all that, the method presented in this work stands out for its simplicity and convenience in implementation, as it does not require any additional complex code designs for extracting MFs from triangulated isosurfaces. This is undeniably one of the main merits of our method.
@@ Acknowledgments
Y.L. would like to thank Cheng Zhao, Charling Tao, Yuting Wang, Shuo Yuan, Zhao Chen, Zhengxiang Li, and Dandan Xu for useful communications. We also thank the anonymous referee for their very detailed and helpful comments, which helped make the paper more informative. This work was supported by National Science Foundation of China (Nos. 12303005, 12273015, 12273020, 11621303, 12173030), National Key R&D Program of China (Nos. 2023YFA1605601, 2023YFA1607800, 2023YFA1607802), and the science research grants from China Manned Space Project with No. CMS-CSST-2021-A03. Y.L. acknowledges support from Shuimu Tsinghua Scholar Program (No. 2022SM173). Y.Y. acknowledges the sponsorship from Yangyang Development Fund. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.
@@ Appendix A The Determinations of Smoothing Lengths
In this appendix, we illustrate our strategy for determining the smoothing lengths adopted in Section 5.2 and display the effects of shot noise on CIC+GS MFs. In previous works (e.g., Park et al. 2005; Choi et al. 2010; Parihar et al. 2014; Kim et al. 2014), smoothing lengths were empirically determined with the restriction $R_{G}\geq\bar{d}\sim\bar{n}_{\mathrm{tracer}}^{-1/3}$, where $\bar{n}_{\mathrm{tracer}}$ is tracer number density. Actually, in our tests, setting $R_{G}=\bar{d}$ cannot provide adequate smoothing for completely eliminating shot noise effects. Our strategy is to strike a balance between removing shot noise effects and retaining sufficient information. To achieve this, $R_{G}$ is determined through an iterative and stepwise refined process, such that the differences of MFs between downgraded field and noise-free field are controlled within $1$-$\sigma$ deviation. We note that this strategy seems to have never been proposed before. Therefore, it is worth exploring it theoretically.
@@ Appendix B The RSD Signals and S/N Ratios Extracted from DTFE MFs
Figure 14: The MFs of subfields. The upper panels are for CIC+MFs, and the lower panels are for DTFE MFs. From left to right, each panel shows the case of halo number density of $1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, $1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$, and $1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$, respectively. For each panel, we show 100 sets of MFs, which are randomly selected from the MFs of 512 subfields. Here, each $V_{i}$ has $n_{\mathrm{b}}=64$ data points, and the unshielded regions indicate the density threshold ranges actually used in our calculations for Fisher forecasts (cf. Table 1).
@@ Appendix C The Methodology of Fisher Forecasts for Errors on Structure Growth Rate
In our scenario, data vectors refer to the MFs in redshift space $V_{i}^{(\mathrm{z})}(\nu)$, and data realizations refer to the $V_{i}^{(\mathrm{z})}(\nu)$ measured from $n_{\mathrm{f}}=8^{3}=512$ subfields (cf. Section 2). Specifically, the covariance matrices are first calculated by using the $512$ subfields with RSDs and then are rescaled by $1/n_{\mathrm{f}}$ [^14] This method is based on an implicit assumption that all subfields are statistically independent. This is not really satisfied in practice. Therefore, it only gives rough estimations (cf. Lacasa & Kunz 2017; Jiang et al. 2023). Since we mainly focus on comparisons of the two schemes for measuring MFs, this issue, in principle, does not affect our conclusions. To solve this issue, one can use a large number of simulation realizations to give accurate but expensive covariance estimates. We leave this for future works.. To calculate the derivatives in Equation C1, we construct two additional halo fields with modified RSDs by artificially increasing and decreasing the peculiar velocity of each halo by a factor of $0.03$, respectively. Then, the response functions are given as $\Delta\boldsymbol{\mu}=\boldsymbol{\mu}_{\mathrm{v}_{+}}-\boldsymbol{\mu}_{\mathrm{v}_{-}}=\Delta V_{i}^{(\mathrm{z})}(\nu)$, where subscripts $\mathrm{v}_{+}$ and $\mathrm{v}_{-}$ indicate the two cases of increasing and decreasing halo peculiar velocities, respectively. Moreover, in RSD studies, there is an approximate relational expression between structure growth rate $f\sigma_{8}(z)$ and halo/galaxy velocity bias $b_{\mathrm{v}}$ (cf. Chen et al. 2018),
@@
In our calculation, we select $n_{\mathrm{b}}$ data points for each $V_{i}(\nu)$. These data points are evenly distributed within $[\nu_{\mathrm{min}},\nu_{\mathrm{max}}]$ in logarithmic coordinate. To avoid any problems sourced from the irreversibility of covariance matrices, the used density threshold range $[\nu_{\mathrm{lower}},\nu_{\mathrm{upper}}]$ should be narrowed as halo number density decreases. This will result in a decrease in the number of data points actually used in calculations. We illustrate this in Fig. 14151515We only display the results of $n_{\mathrm{b}}=64$, considering the striking similarity between these results and those of $n_{\mathrm{b}}=32$. (cf. the unshielded regions) and Table 1. In order to draw conclusions as reliable as possible, two $n_{\mathrm{b}}$ cases are tested, namely $n_{\mathrm{b}}=32$ (cf. Fig. 11) and $n_{\mathrm{b}}=64$ (cf. Fig. 15). This provides a consistency check for our methodology. In general, in comparison with Fig. 11, the predictive errors in Fig. 15 are relatively smaller, depending on different cases. Specifically, as the number of data points increases, we find the followings:
@@
    For DTFE MFs, the predictive errors remain relatively stable in the case of $n_{\mathrm{h}}=1.6\times 10^{-2}\,(h^{-1}\mathrm{Mpc})^{-3}$, but show a noticeable decrease in the cases of $n_{\mathrm{h}}=1.6\times 10^{-3}\,(h^{-1}\mathrm{Mpc})^{-3}$ and $n_{\mathrm{h}}=1.6\times 10^{-4}\,(h^{-1}\mathrm{Mpc})^{-3}$.